OpenStudy (anonymous):

-n - 24 = 5 - explain please

Directrix (directrix):

Is the task to solve for n?

OpenStudy (anonymous):

-n-24 = 5 +24 +24 ------------- -n = 29 *-1 *-1 ------------- n = -29 what you do to one side of the equation, you MUST do to the other side. Your goal is also to isolate the variable.

Directrix (directrix):

@chandhooguy OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

thank you chandhoogy i always get confused around that part.

OpenStudy (anonymous):

No problem. If I helped you, it would help me if you gave me a medal. Thanks!